OpenStudy (anonymous):

Verify the identity. cos (x - y) - cos (x + y) = 2 sin x sin y

OpenStudy (anonymous):

I can't really figure out what identities to use

OpenStudy (anonymous):

Cos(x - y) = Cos(x)*Cos(y) + Sin(x)*Sin(y) Cos(x + y) = Cos(x)*Cos(y) - Sin(x)*Sin(y)

OpenStudy (anonymous):

ah I see! I figured it out! thank you!

OpenStudy (anonymous):

npz